How many grams are there in 1.56 moles of potassium sulfate


1
Expert's answer
2021-04-09T01:46:23-0400

molar mass of "K_2SO_4=39\times 2+32+16\times4=174g"

mass of "K_2SO_4=no. of \space moles \times molar \space mass=1.56\times 174g=271.44g"